angelcook2007:

The height of water shooting from a fountain is modeled by the function f(x) = −4x2 + 24x − 29 where x is the distance from the spout in feet. Complete the square to determine the maximum height of the path of the water.

surjithayer:

do you know how to solve it?

angelcook2007:

no

surjithayer:

\[f(x)=-4x^2+24x-29\] \[=-4(x^2-6x+...)-29\] can you complete the square?

surjithayer:

\[=-4[x^2-6x+{(coefficient~of~x)/2)}^2-((coefficient ~of~x)/2)^2]-29\]

angelcook2007:

I got it, thanks mate

surjithayer:

\[=-4(x^2-6x+(-6/2)^2-(-6/2)^2)-29\] \[=-4(x-3)^2-4(-9)-29\] \[=-4(x-3)^2+36-29\] \[=-4(x-3)^2+7\] so maximum height of water=7 ft.

angelcook2007:

Thank you, I ended up making a 100